# Heterogeneous Mixed Traffic Control and Coordination

### v1

@@ IV-C Traffic Heterogeneity Evaluation
The incorporation of RVs into traffic systems poses unique challenges when considering the diverse range of vehicle types present on urban roadways. These challenges arise from varying vehicle sizes, speeds, and maneuvering capabilities, which complicate traffic flow coordination including intersection management. Here we examine how different vehicle types—passenger cars, pickups, vans, semi-trailers, and trucks—respond to increasing RV rates in terms of traffic efficiency and environmental impact.
@@ IV-C1 Average Waiting Time
Fig. 3 (top row) presents the average waiting times at intersections for different vehicle types across RV rates, compared to HV-Signalized. We observe that all vehicle types experience a decrease in waiting time as RV rate increases, indicating a broad positive impact of RV introduction on traffic efficiency. However, the magnitude of improvement varies across vehicle types.
Figure 3: Average Waiting Time (top row), CO2 Emissions (middle row), and Fuel Consumption (bottom row) for different vehicle types across RV rates (10% to 90%) compared to HV-Signalized. As RV rate increases, waiting times generally decrease for all vehicle types, outperforming HV-Signalized starting from 40% RV penetration. CO2 emissions and fuel consumption show slight increases but remain below HV-Signalized levels for all RV scenarios. Larger vehicles like trucks and semi-trailers show more dramatic increases in emissions and fuel consumption as RV rates rise, due to their size and weight. In contrast, passenger cars and other smaller vehicles exhibit moderate increases in environmental impact while still benefiting from significant waiting time reductions, demonstrating a balanced improvement across all metrics.
@@
TABLE IV: Percentage improvements in waiting times for vehicle types at different RV rates compared to HV-Signalized. Bold values indicate the highest improvement for each vehicle type, all occurring at 90% RV rate. This trend demonstrates rarity advantage for less common vehicles and the benefit of a more homogeneous environment for passenger cars. Trucks show the most significant improvement (87.03%), followed by pickups (77.52%). Passenger cars (both RV and non-RV), despite being the most common, still achieve substantial improvements (52.46%) due to the reduced heterogeneity of the traffic environment at high RV rates.
@@
To understand this trend better, we present the percentage improvements in waiting times for the vehicle types at key RV rates compared to HV-Signalized in Table IV. The disproportionate benefit for less-frequent vehicles like trucks and pickups can be attributed to their ability to take advantage of the gaps created by RV coordination without contributing significantly to overall congestion. Their infrequent presence at intersections means they are more likely to encounter optimized traffic conditions created by the RVs. Similarly, semi-trailers’ moderate improvement compared to trucks might be due to their higher frequency in traffic, leading to more vehicle interactions and potentially more complex maneuvering requirements. In contrast, more common vehicles like passenger cars and vans see moderate improvements, likely due to their frequent interactions with other vehicles.
@@ IV-C2 CO2 Emissions and Fuel Consumption
We present the trends in CO2 emissions and fuel consumption across different vehicle types as RV rate increases in Fig. 3 (middle and bottom rows).
For CO2 emissions, we observe a general increasing trend for all vehicle types as RV rates rise. Passenger cars show a moderate increase, with emissions rising from 2,333 mg/s at 10% RV rate to 2,917 mg/s at 90% RV rate, a 25% increase. Pickups exhibit a similar trend, experiencing a 24% rise, with emissions increasing from 2,261 mg/s to 2,813 mg/s. Larger vehicles demonstrate more substantial increases in CO2 emissions. Trucks, for example, see their emissions rise from 8,497 mg/s at 10% RV rate to 12,946 mg/s at 90% RV rate, leading to a 52% increase. Semi-trailers show the most dramatic change, with emissions going from 7,554 mg/s to 12,165 mg/s, a significant 61% increase. The overall trend shows that CO2 emissions rise more sharply for larger vehicles as RV rate increases.
Fuel consumption patterns closely mirror the CO2 emission trends. Semi-trailers show the most significant increase, with fuel consumption rising from about 3.2 ml/s at 10% RV rate to 5.2 ml/s at 90% RV rate, a 62.5% increase. Trucks follow a similar pattern, with consumption increasing from 3.6 ml/s to 5.6 ml/s, a 55.5% rise. Passenger cars, pickups, and vans exhibit lower overall fuel consumption but still show an increasing trend with higher RV rates. For instance, passenger car fuel consumption rises from 1.0 ml/s to 1.25 ml/s, a 25% increase as RV rate goes from 10% to 90%. Overall, larger vehicles experience the greatest rise in fuel consumption at higher RV rates, while smaller vehicles see more moderate increases.
@@ IV-C3 Space Headway
As shown in Fig. 4, we observe a consistent reduction in space headways across all vehicle types as RV rate increases. The impact of RVs on space headways varies across vehicle types. Passenger cars, which dominate the traffic composition, show a marked decrease in space headways from 57.49 m at 10% RV rate to 35.78 m at 90% RV rate, a reduction of about 38%. Larger vehicles such as semi-trailers, which typically require more space due to their size and slower acceleration, also exhibit reduced space headways, with space headways decreasing from 70.55 m to 53.21 m (25% reduction) as RV rate increases from 10% to 90%. These results underscore the potential of RVs in facilitating streamlined traffic flows.
Figure 4: Space headways for different vehicle types at varying RV penetration rates. As RV rate increases, space headways decrease across all vehicle types, leading to better road space utilization. Passenger cars show the largest reduction. Trucks and semi-trailers see smaller reductions, with a plateau effect at higher RV rates.
In case of less occurring vehicles (e.g., trucks and pickups) at higher RV rates, we observe a plateau effect, where headways stabilize or slightly increase. This could be attributed to the fact that low-frequency vehicles interact less frequently at intersections and face fewer congestion issues in high-RV scenarios. Therefore, they do not experience the same pressure to reduce headways as passenger cars do. The changing vehicle type mix at higher RV rates may also contribute to this effect.
@@ V CONCLUSION AND FUTURE WORK
This study investigates the impact of heterogeneous mixed traffic control using RVs in complex environments. Our results show that higher RV rates significantly reduce average waiting times for all vehicle types, with reductions of up to 86% and 91% compared to HVs at signalized and unsignalized intersections, respectively. Notably, less common vehicles like trucks benefit the most from RV coordination. While RVs lead to increased emissions and fuel consumption due to higher speeds, the overall efficiency gains surpass those of signalized intersections. Additionally, we observe consistent reductions in space headways across all vehicle types, with the largest decrease for passenger cars.
There exist several future directions. First, we would like to expand the testing of heterogeneous mixed traffic to larger-scale networks by leveraging existing traffic simulation, reconstruction, and prediction techniques [40, 41, 42, 43, 44, 45, 46, 47]. Second, we plan to enhance the capabilities of RVs by integrating advanced autonomous features in both 2D simulation [48, 49, 50] and 3D simulation [51, 52]. Finally, testing communication strategies among RVs under both short and long communication ranges through network optimization [53] could further improve the applicability of our approach.

### v2

@@ IV-C Traffic Heterogeneity Evaluation
The integration of RVs into traffic systems presents challenges due to the diverse vehicle types on urban roads. Differences in size, speed, and maneuverability complicate coordination, particularly at intersections. This study analyzes how various vehicle types—passenger cars, pickups, vans, semi-trailers, and trucks—respond to rising RV rates in terms of traffic efficiency and environmental impact.
@@ IV-C1 Average Waiting Time
Fig. 3 (top row) presents the average waiting times at intersections for different vehicle types across RV rates, compared to HV-Signalized. As the RV rate approaches 0% in our unsignalized intersections, the results align with the HV-Unsignalized baseline. This explains why the waiting times at 10% RV rate are higher than HV-Signalized baseline for all vehicle types. As demonstrated in our overall results, the HV-Unsignalized scenario performs significantly worse than HV-Signalized on average. We observe that all vehicle types experience a decrease in waiting time as RV rate increases, indicating a broad positive impact of RV introduction on traffic efficiency. However, the magnitude of improvement varies across vehicle types.
Figure 3: Waiting Time (top row), CO2 Emissions and Fuel Consumption (bottom row) for different vehicle types across RV rates (10% to 90%) compared to HV-Signalized, averaged over a 1000-second run. As RV rate increases, waiting times generally decrease for all vehicle types, outperforming HV-Signalized starting from 40% RV penetration. Due to their near-identical trends in SUMO’s emission model, a single line is used to represent both CO2 emissions and fuel consumption in the bottom row. These metrics slightly increase but stay below HV-Signalized levels. Larger vehicles, like trucks and semi-trailers, show more pronounced increases due to size and weight, while smaller vehicles benefit from reduced waiting times with moderate environmental impact.
@@
TABLE IV: Percentage improvements in waiting times for vehicle types compared to HV-Signalized. Bold values indicate highest improvement for each vehicle type at 90% RV rate, demonstrating rarity advantage for less common vehicles and benefits of reduced heterogeneity for passenger cars.
@@
To better understand this trend, Table IV shows percentage improvements in waiting times for each vehicle type at key RV rates compared to HV-Signalized. The observed “rarity advantage,” where less frequent vehicles like trucks and pickups show greater improvements than common vehicles, can be attributed to temporal spacing, i.e., their ability to take advantage of the gaps created by RV coordination without contributing significantly to overall congestion. Their infrequent presence at intersections means they are more likely to encounter optimized traffic conditions created by the RVs. Similarly, semi-trailers’ moderate improvement compared to trucks might be due to their higher frequency in traffic, leading to more vehicle interactions and potentially more complex maneuvering requirements. In contrast, passenger cars (70% of traffic) frequently encounter other vehicles competing for intersection access. Interestingly, we observe a slight increase in waiting times for most vehicle types when the RV rate reaches 60–70%, before decreasing again at higher rates. This non-monotonic pattern suggests a potential phase transition of coordination complexity within traffic dynamics, as the system shifts from HV-dominated to RV-dominated. At higher RV rates (80–90%), the greater uniformity in vehicle control strategies allows for more efficient system-wide coordination, resulting in the resumed decrease in waiting times. The dip may also stem from a cascading effect in the sequential intersection layout, where efficient RV flow from upstream intersections leads to dense platoons that challenge coordination downstream. This phenomenon shows the complex, non-linear relationship between RV rates and traffic efficiency under heterogeneous settings.
@@ IV-C2 CO2 Emissions and Fuel Consumption
Fig. 3 (bottom row) shows CO2 emissions and fuel consumption trends across vehicle types as RV rate increases. For CO2 emissions, we observe a general increasing trend for all vehicle types as RV rates rise. Passenger cars show a moderate increase, with emissions rising from 2,333 mg/s at 10% RV rate to 2,917 mg/s at 90% RV rate, a 25% increase. Pickups exhibit a similar trend, experiencing a 24% rise. Larger vehicles demonstrate more substantial increases in CO2 emissions. Trucks, For example, trucks see their emissions rise from 8,497 mg/s at 10% RV rate to 12,946 mg/s at 90% RV rate, leading to a 52% increase. Semi-trailers show the most dramatic change, with emissions going from 7,554 mg/s to 12,165 mg/s, a significant 61% increase. The overall trend highlights that CO2 emissions rise more sharply for larger vehicles as RV rate increases. Fuel consumption patterns closely mirror the CO2 emission trends. Semi-trailers show the most significant increase, with fuel consumption rising from about 3.2 ml/s at 10% RV rate to 5.2 ml/s at 90% RV rate, a 62.5% increase. Trucks follow a similar pattern, with consumption increasing from 3.6 ml/s to 5.6 ml/s, a 55.5% rise. Passenger cars, pickups, and vans exhibit lower overall fuel consumption but still show an increasing trend with higher RV rates. For instance, passenger car fuel consumption rises from 1.0 ml/s to 1.25 ml/s, a 25% increase as RV rate goes from 10% to 90%. Overall, larger vehicles experience the greatest rise in fuel consumption at higher RV rates, while smaller vehicles see more moderate increases.
@@ IV-C3 Space Headway
As shown in Fig. 4, we observe a consistent reduction in space headways across all vehicle types as RV rate increases. The impact of RVs on space headways varies across vehicle types. Passenger cars, which dominate the traffic composition, show a marked decrease in space headways from 57.49 m at 10% RV rate to 35.78 m at 90% RV rate, a reduction of about 38%. Larger vehicles such as semi-trailers, which typically require more space due to their size and slower acceleration, also exhibit reduced space headways, with space headways decreasing from 70.55 m to 53.21 m (25% reduction) as RV rate increases from 10% to 90%. These results underscore the potential of RVs in facilitating streamlined traffic flows. In case of less occurring vehicles (e.g., trucks and pickups) at higher RV rates, we observe a plateau effect, where headways stabilize or slightly increase. This could be attributed to the fact that low-frequency vehicles interact less frequently at intersections and face fewer congestion issues in high-RV scenarios. Therefore, they do not experience the same pressure to reduce headways as passenger cars do. The changing vehicle type mix at higher RV rates may also contribute to this effect.
Figure 4: Average space headways for different vehicle types at varying RV penetration rates. As RV rate increases, space headways decrease across all vehicle types, leading to better road space utilization. Passenger cars show the largest reduction. Trucks and semi-trailers see smaller reductions, with a plateau effect at higher RV rates.
@@ V CONCLUSION AND FUTURE WORK
This study investigates heterogeneous mixed traffic control using RVs in complex environments. Results show that higher RV rates significantly reduce average waiting times for all vehicle types—up to 86% and 91% lower than HVs at signalized and unsignalized intersections, respectively. The rarity advantage suggests certain vehicle types benefit more from RV control, indicating that adaptive strategies with vehicle-type awareness could further boost efficiency. While RVs raise emissions and fuel use due to higher speeds, overall efficiency still surpasses signalized baselines. We also observe reduced space headways across all vehicle types, with the largest drop for passenger cars.
There exist several future directions. First, testing RV operation in more unseen environments—such as roundabouts, three-way intersections, and large-scale networks—and with imperfect information from local sensors or lossy V2V communication would better assess policy robustness and generalization. Second, adopting a continuous action space could enable more efficient coordination than the current Stop/Go model. Third, this study uses only passenger cars as RVs for tractability; extending to other vehicle types may allow vehicle-specific reward shaping (e.g., fuel-efficient truck behaviors). Finally, broader comparisons with alternative RL or optimization-based intersection management strategies remain an important future direction.
Acknowledgments
This research is supported by NSF 2153426, 2332210, 2524239, and 2129003. The authors also thank NVIDIA and the Tickle College of Engineering at University of Tennessee, Knoxville for their support.